# Un-mixing Test-time Adaptation under Heterogeneous Data Streams

### v2

@@ VII-B1 FreDA improves across diverse distribution shifts
Our method consistently attains the lowest error rates across all evaluated datasets (see TABLE V and III). Notably, on the Camelyon17 dataset, FreDA reduced the error rate to 27.9%, outperforming the next best method by 5.9%. This significant improvement is particularly noTABLE where other approaches falter if compared with no training (TBN), meaning they struggle to adapt to the complex medical imaging data. By effectively handling high variability and intricate patterns in the data, FreDA maintains superior accuracy and adaptability.
@@ VII-B2 FreDA remains stable under various batch size
To simulate deployment with constrained batch sizes, we evaluate models under both varying batch sizes and mixed distributions. In Figure 5, we present the results on CIFAR-10-C, CIFAR-100-C, and ImageNet-C using batch sizes ranging from 200 (64) down to 1. Unlike other methods that significantly degrade as batch size decreases – for example the error rate of DeYO increases from 27.7% to 89.8% when batch size drops from 200 to 1 on CIFAR-10-C – FreDA consistently maintains strong performance. This stability demonstrates FreDA’s robustness, making it highly suiTABLE for real-world applications where large batches is not always feasible.
TABLE IV: Ablation study of FreDA.
DT SS SA C10 C100 IN
44.1 82.5 86.4
24.8 54.2 81.2
29.6 37.5 71.0
39.4 71.7 92.9
24.3 36.3 69.4
27.7 36.2 65.9
24.4 50.2 77.7
22.9 34.7 67.2
VII-B3 FreDA enhances adaptation via synergistic designs
This section validates our designs by ablating three key modules – Decentralized Training (DT), Sample Selection (SS), and Sample Augmentation (SA). The baseline here leverages only the entropy loss. From TABLE VI, we have the following observations: 1) Implementing decentralized training and sample selection alone results in substantial improvements, reducing error rates dramatically across all datasets. 2) Sample augmentation alone has the possibility to increase error rates, suggesting that although this approach introduces useful variability, it may introduce unexpected noise under the absence of proper selection or decentralized training. 3) The combined approach delivers the best performance across all datasets, showing the synergistic effect of our different designs.
TABLE V: Classification error rate ($\downarrow$) on ImageNet-C under Mixed Distribution Shifts using ViT-Base backbone.
Baseline & Methods Gauss. Shot Impul. Defoc. Glass Motion Zoom Snow Frost Fog Brig. Contr. Elast. Pixel JPEG Avg.
@@ VII-B3 FreDA enhances adaptation via synergistic designs
TENT (ICLR 21’) 60.6 60.4 59.6 63.6 67.8 57.1 61.2 55.0 48.8 47.4 28.6 66.7 53.9 50.4 44.4 55.0
ETA (ICML 22’) 59.3 57.8 57.9 58.8 62.8 52.5 58.2 51.0 46.4 44.2 28.8 58.3 51.1 46.9 41.9 51.7
AdaContrast (CVPR 22’) 64.8 63.4 63.3 72.8 76.6 73.7 74.6 67.7 48.0 89.6 30.2 93.2 60.8 57.3 46.3 65.5
CoTTA (CVPR 22’) 89.4 92.0 88.9 93.6 92.6 90.6 86.5 94.9 88.2 86.6 75.8 96.5 85.7 93.5 84.6 89.3
SAR (ICLR 23’) 58.9 57.6 57.6 59.4 63.6 53.0 58.5 52.3 47.1 45.4 28.3 61.6 51.4 47.4 42.0 52.3
RoTTA (CVPR 23’) 64.4 65.6 63.7 67.6 71.3 59.8 64.1 52.7 43.5 48.6 27.9 78.5 54.3 60.4 50.1 58.2
RDumb (NeurIPS 23’) 59.7 58.5 58.5 60.0 64.1 54.0 59.0 52.0 46.7 44.5 28.6 61.2 51.9 48.3 42.6 52.6
DeYO (ICLR 24’) 60.0 58.6 58.8 58.8 62.4 61.9 50.9 46.7 51.9 45.2 29.7 55.7 51.6 45.8 42.8 52.1
@@
VII-C More analysis and Discussion
VII-C1 Performance with Transformer Backbone
In addition to evaluating our model on commonly compared CNN backbones, we further assess its performance under transformer-based architecture, specifically using the ViT-Base backbone. Here, we report results on the ImageNet-C benchmark using ViTBase-LN [45] (see TABLE V), where the pretrained weights are obtained from torchvision. Importantly, all experimental configurations are kept consistent with those used in the CNN-based experiments. As shown, our method continues to deliver strong performance, demonstrating its robustness and adaptability across different backbone architectures.
TABLE VI: Sensitivity analysis on different datasets.
CLUSTER_NUM 2 4 8 16
CIFAR10-C 23.0 22.9 23.2 24.7
CIFAR100-C 34.8 34.7 34.7 35.6
IN-C (ResNet) 68.6 67.2 67.1 70.5
IN-C (ViT) 50.3 48.8 49.9 50.0
KMEANS_SIZE 256 512 1024 2048
CIFAR10-C 23.0 22.9 23.0 22.9
CIFAR100-C 34.6 34.7 34.8 34.8
IN-C (ResNet) 69.0 67.2 67.6 67.0
IN-C (ViT) 49.0 48.8 48.7 48.8
COMM_INTERVAL 1 10 100 1000
CIFAR10-C 22.6 22.9 22.6 22.0
CIFAR100-C 34.7 34.7 34.9 43.2
IN-C (ResNet) 67.1 67.2 67.2 67.4
IN-C (ViT) 48.4 48.8 48.8 48.7
PERT_MAGNITUDE 0.0 0.1 0.2 0.3
CIFAR10-C 24.3 22.9 22.5 22.2
CIFAR100-C 36.3 34.7 34.9 34.9
IN-C (ResNet) 69.3 67.2 67.0 66.9
IN-C (ViT) 49.6 48.8 48.5 48.6
VII-C2 Sensitivity Study
We investigate the impact of the key hyperparameters: CLUSTER_NUM, KMEANS_SIZE, COMM_INTERVAL, and PERT_MAGNITUDE. From TABLE VI, we have the following observations: 1) The choice of CLUSTER_NUM influences model performance, especially on more complex datasets. While performance remains relatively stable even with a small number of clusters (e.g., two clusters), increasing the number of clusters beyond four tends to lead to slight performance degradation, particularly on datasets like ImageNet-C. This suggests that while more clusters can improve adaptation capacity, excessively increasing the number can introduce overfitting and reduce generalization. Therefore, a moderate number of clusters, around 4, appears to strike a good balance between adaptation flexibility and maintaining model robustness. 2) Varying KMEANS_SIZE from 256 to 2048 results in stable performance across all datasets, indicating that our method is robust to changes in cluster sizes. 3) Our approach shows general robustness to communication frequency (varying COMM_INTERVAL from 1 to 1,000). On simpler datasets such as CIFAR10-C, infrequent communication (e.g., interval $f=1000$) performs best, likely due to effective independent learning. In contrast, for complex datasets, more frequent communication (e.g., $f=1$) improves performance, likely by mitigating divergence among local branches and ensuring model consistency. 4) Adjusting the perturbation magnitude (PERT_MAGNITUDE) has a noticeable effect, particularly on simpler datasets. Increasing perturbation improves performance, suggesting that spectral augmentation contributes to better generalization. However, for more complex datasets such as CIFAR100-C, higher perturbation levels lead to slight decline in performance, likely due to noise disrupting the alignment of finer features. A perturbation magnitude of 0.1 appears to strike an effective balance, offering the benefits of augmentation without negatively impacting feature alignment.
TABLE VII: Classification error rate ($\downarrow$) on CIFAR-10-C (C10), CIFAR-100-C (C100), and ImageNet-C (IN) using ResNet-50 & ViT-Base backbones under Continual Setting, averaged over 15 corruptions.
@@ VII-C2 Sensitivity Study
TENT (ICLR 21’) 20.0 62.2 62.6 54.5
ETA (ICML 22’) 17.9 32.2 60.2 49.8
AdaContrast (CVPR 22’) 18.5 33.5 65.5 57.0
CoTTA (CVPR 22’) 16.5 32.8 63.1 77.0
SAR (ICLR 23’) 20.4 32.0 61.9 51.7
RoTTA (CVPR 23’) 19.3 34.8 67.3 58.3
RDumb (NeurIPS 23’) 17.8 34.1 90.6 50.2
DeYO (ICLR 24’) 87.0 98.1 90.6 94.3
UnMix-TNS (ICLR 24’) 24.9 32.7 75.4 -
@@
VII-C3 Performance under continual settings
@@
VIII Conclusion and Future Work
This paper advances Test-Time Adaptation (TTA) by tackling the real-world complexities of heterogeneous data streams. Our decentralized approach precisely manages diverse data shifts, improving model adaptation in varied settings. By integrating Fourier-based augmentation, we expand the range of confident samples for each distribution shift, further boosting model performance. The experimental results underscore the efficacy of FreDA, highlighting its potential to influence the field and guide future research in adapting to dynamic and diverse data shifts. While FreDA addresses a critical challenge in handling heterogeneous data streams, providing a solid pipeline for this issue, there are still avenues for further enhancement. Our current aggregation approach, which averages models based on cluster counts, has been effective in solving the problem at hand. However, exploring alternative strategies – such as weighting models by the divergence between clusters – might lead to incremental improvements. Additionally, refining the sample selection process from a original sample-level focus to a more granular patch-level could extend FreDA’s applicability to tasks such as segmentation, further enhancing its versatility in real-world scenarios.
-A Adaptation Scenarios
-A1 Mixed Domains
: In this scenario, the model processes a long sequence of test samples where each sample $x_{i}\sim p_{e_{i}}(x)$ is independently drawn from a randomly selected target domain $\mathcal{D}_{e_{i}}\in\{\mathcal{D}_{t_{1}},\mathcal{D}_{t_{2}},\dots,\mathcal{D}_{t_{N}}\}$ and a randomly selected class $c_{i}$ among classes $\{1,2,\dots,C\}$. The sequence is represented as:
@@ VIII Conclusion and Future Work
-A2 Continual Domain Adaptation
: In this setting, the model adapts online to a sequentially presented series of test domains, where each domain shift occurs only once and does not reappear. The sequence progresses through distinct target domains in a fixed order $D_{1}\to D_{2}\to\dots\to D_{N}$, with samples from each domain appearing contiguously. The class labels within each domain are independently and randomly selected. The sequence is structured as:
$\begin{split}&\underbrace{\{(x_{1}^{D_{1},c_{1}}),(x_{2}^{D_{1},c_{2}}),\dots,(x_{l_{1}}^{D_{1},c_{l_{1}}})\}}_{\text{Samples from domain }D_{1}}\\ &\to\underbrace{\{(x_{l_{1}+1}^{D_{2},c_{l_{1}+1}}),\dots,(x_{l_{1}+l_{2}}^{D_{2},c_{l_{1}+l_{2}}})\}}_{\text{Samples from domain }D_{2}}\\ &\to\dots\to\underbrace{\{(x_{l_{1}+\dots+l_{N-1}+1}^{D_{N},c_{l_{1}+\dots+l_{N-1}+1}}),\dots,(x_{k}^{D_{N},c_{k}})\}}_{\text{Samples from domain }D_{N}}\end{split}$
@@
Figure 6: Examples from ImageNet-C under common image corruptions. The images showcase a range of corruption types (e.g., noise, blur, and weather distortions) at varying severity levels.
@@
-B Dataset Visualization
@@
-C Relation to Frequency Domain Learning
Recent advances highlight frequency-based techniques as powerful tools for domain transfer. In domain generalization, frequency analysis has revealed critical insights into model robustness and learning dynamics [46, 47, 48, 49, 50, 51, 52, 53]. For domain adaptation, interpolating image amplitude spectra across styles has proven effective in reducing domain gaps and preventing overfitting to low-level statistics [54, 55, 56, 57]. Motivated by these advancements, we discover that frequency information inherently captures domain characteristics, making it a valuable medium for decoupling mixed target domains – an aspect largely unexplored in prior work. On this basis, we propose FreDA, the first framework addressing TTA under mixed domain shifts in Fourier space, with a decentralized adaptation and perturbation mechanism.
-D Relation to Multi-Target Domain Adaptation
TTA under mixed distribution shifts shares similarities with multi-target unsupervised domain adaptation (MT-UDA) [11, 12, 58, 59], yet diverges critically in complexity: while MT-UDA assumes static, predefined target domains and leverages labeled source data for explicit domain alignment, TTA operates with no access to source data and must adapt to dynamic, unpredicTABLE target streams in an online manner. This eliminates direct source-target discrepancy computation and demand robust incremental adaptation rather than offline multi-domain optimization. These differences highlight the need for novel methodologies tailored specifically to TTA, going beyond the solutions developed for MT-UDA.
-E Relation to Decentralized, Federated, Distributed Learning
This work also intersects with decentralized, federated, and distributed learning due to splitting data batches into disjoint subsets and applying decentralized model adaptation. First, while decentralized learning focuses on non-i.i.d. data that is naturally distributed across multiple nodes [60], our approach starts with a centralized batch of target samples. By proactively splitting this data into disjoint subsets, we expose latent non-i.i.d. characteristics, enabling the effective use of decentralized learning techniques. Second, federated learning considers data privacy and model collaborations within decentralized learning [61]. In our case, as target samples are mixed in a batch, data privacy is not a concern. However, similar to federated learning, our approach also involves weight aggregation over subnetworks to enhance their base models before next batch adaptation. Third, distributed learning aims to improve training efficiency on large-scale datasets by partitioning data for synchronized training [62]. In contrast, our method operates in a real-time fine-tuning context with limited data at one time, hence scalability is less of a concern.

### v4

@@ VII-B1 FreDA improves across diverse distribution shifts
Our method consistently attains the lowest error rates across all evaluated datasets (see TABLE II and III). Notably, on the Camelyon17 dataset, FreDA reduced the error rate to 27.9%, outperforming the next best method by 5.9%. This significant improvement is particularly notable where other approaches falter if compared with no training (TBN), meaning they struggle to adapt to the complex medical imaging data. By effectively handling high variability and intricate patterns in the data, FreDA maintains superior accuracy and adaptability.
@@ VII-B2 FreDA remains stable under various batch size
To simulate deployment with constrained batch sizes, we evaluate models under both varying batch sizes and mixed distributions. In Fig. 5, we present the results on CIFAR-10-C and CIFAR-100-C using batch sizes ranging from 200 down to 1. Unlike other methods that significantly degrade as batch size decreases – for example the error rate of DeYO increases from 27.7% to 89.8% when batch size drops from 200 to 1 on CIFAR-10-C – FreDA consistently maintains strong performance. This stability demonstrates FreDA’s robustness, making it highly suitable for real-world applications where large batches is not always feasible.
VII-B3 FreDA enhances adaptation via frequency-based designs
This section validates our designs by ablating two key modules – Frequency-based Decentralized learning (FD) and Frequency-based Augmentation (FA) – from the full model. It is worth mentioning that our FreDA is built upon TENT (our baseline) by integrating both FD and FA. From Fig. 6, we observe the following: a) Removing FD leads to a clear performance drop, underscoring the importance of our proposed decentralized adaptation paradigm in handling mixed distribution shifts. b) Removing FA results in a surprisingly large error rate increase on CIFAR-100-C. In more challenging classification scenarios (e.g., CIFAR-100-C vs. CIFAR-10-C), the target domain typically exhibits sparser class-wise coverage. Augmentation enriches the target sample diversity while preserving semantic consistency, thereby mitigating overfitting to spurious patterns and enhancing robustness under severe and complex distribution shifts. Nevertheless, even without FA, our method still significantly outperforms the baseline model, further confirming the critical role of our decentralized adaptation framework.
Figure 6: Classification error rate (↓) on CIFAR-10-C/100-C with various batch size under Mixed Distribution Shifts.
TABLE IV: Classification error rate ($\downarrow$) on ImageNet-C under Mixed Distribution Shifts using ViT-Base backbone.
Methods Gauss. Shot Impul. Defoc. Glass Motion Zoom Snow Frost Fog Brig. Contr. Elast. Pixel JPEG Avg.
@@ VII-B3 FreDA enhances adaptation via frequency-based designs
TENT 60.6 60.4 59.6 63.6 67.8 57.1 61.2 55.0 48.8 47.4 28.6 66.7 53.9 50.4 44.4 55.0
ETA 59.3 57.8 57.9 58.8 62.8 52.5 58.2 51.0 46.4 44.2 28.8 58.3 51.1 46.9 41.9 51.7
AdaContrast 64.8 63.4 63.3 72.8 76.6 73.7 74.6 67.7 48.0 89.6 30.2 93.2 60.8 57.3 46.3 65.5
CoTTA 89.4 92.0 88.9 93.6 92.6 90.6 86.5 94.9 88.2 86.6 75.8 96.5 85.7 93.5 84.6 89.3
SAR 58.9 57.6 57.6 59.4 63.6 53.0 58.5 52.3 47.1 45.4 28.3 61.6 51.4 47.4 42.0 52.3
RoTTA 64.4 65.6 63.7 67.6 71.3 59.8 64.1 52.7 43.5 48.6 27.9 78.5 54.3 60.4 50.1 58.2
RDumb 59.7 58.5 58.5 60.0 64.1 54.0 59.0 52.0 46.7 44.5 28.6 61.2 51.9 48.3 42.6 52.6
DeYO 60.0 58.6 58.8 58.8 62.4 61.9 50.9 46.7 51.9 45.2 29.7 55.7 51.6 45.8 42.8 52.1
@@
Figure 7: Parameter study with respect to CLUSTER_NUM, KMEANS_SIZE, COMM_INTERVAL & PERT_MAGNITUDE (from top to bottom) under Mixed Distribution Shifts.
VII-B4 FreDA adapts robustly to hyperparameter variations
We investigate the impact of the key hyperparameters: CLUSTER_NUM, KMEANS_SIZE, COMM_INTERVAL, and PERT_MAGNITUDE. From Fig. 7, we have the following observations: a) CLUSTER_NUM is largely insensitive in the range of 2–8: performance varies only marginally across CIFAR10/100-C and remains stable on average, with a clear drop emerging only at the extreme setting of 16. We attribute the degradation at 16 to over-partitioning, where the model tends to overfit to subtle, cluster-specific variations rather than capturing domain-invariant patterns. Therefore, a moderate number of clusters, around 4, appears to strike a good balance between adaptation flexibility and maintaining model robustness. b) Varying KMEANS_SIZE from 256 to 2048 results in stable performance across all datasets, indicating that our method is robust to changes in cluster sizes. c) Our approach shows general robustness to communication frequency (varying COMM_INTERVAL from 10 to 1,000). On simpler datasets such as CIFAR10-C, infrequent communication (e.g., interval $f=1000$) performs best, likely due to effective independent learning. In contrast, for complex datasets (e.g. CIFAR100-C), more frequent communication (e.g., $f=10$) improves performance, likely by mitigating divergence among local branches and ensuring model consistency. d) Introducing perturbation (PERT_MAGNITUDE) from 0 to a positive value yields a clear performance gain, confirming the effectiveness of our frequency-based augmentation strategy. Furthermore, when varying the magnitude between 0.1 and 0.3, FreDA maintains robust performance across datasets, indicating its resilience to moderate perturbation levels.
TABLE V: Classification error rate ($\downarrow$) on CIFAR-10-C (C10), CIFAR-100-C (C100), and ImageNet-C (IN) using ResNet-50 & ViT-Base backbones under Continual Setting, averaged over 15 corruptions.
@@ VII-B4 FreDA adapts robustly to hyperparameter variations
TENT 20.0 62.2 62.6 54.5
ETA 17.9 32.2 60.2 49.8
AdaContrast 18.5 33.5 65.5 57.0
CoTTA 16.5 32.8 63.1 77.0
SAR 20.4 32.0 61.9 51.7
RoTTA 19.3 34.8 67.3 58.3
RDumb 17.8 34.1 90.6 50.2
DeYO 87.0 98.1 90.6 94.3
UnMix-TNS 24.9 32.7 75.4 -
@@
VII-C Performance with Transformer Backbone
In addition to evaluating our model on commonly compared CNN backbones, we further assess its performance under transformer-based architecture, specifically using the ViT-Base backbone. Here, we report results on the ImageNet-C benchmark using ViTBase-LN [49] (see TABLE IV), where the pretrained weights are obtained from torchvision. Importantly, all experimental configurations are kept consistent with those used in the CNN-based experiments. As shown, our method continues to deliver strong performance, demonstrating its robustness and adaptability across different backbone architectures.
VII-D Performance under Continual Settings
@@
VIII More Details and Discussion
VIII-A Adaptation Scenarios
VIII-A1 Mixed Domains
In this scenario, the model processes a long sequence of test samples where each sample $x_{i}\sim p_{e_{i}}(x)$ is independently drawn from a randomly selected target domain $\mathcal{D}_{e_{i}}\in\{\mathcal{D}_{t_{1}},\mathcal{D}_{t_{2}},\dots,\mathcal{D}_{t_{N}}\}$ and a randomly selected class $c_{i}$ among classes $\{1,2,\dots,C\}$. The sequence is represented as:
@@ VIII-A1 Mixed Domains
VIII-A2 Continual Domain Adaptation
In this setting, the model adapts online to a sequentially presented series of test domains, where each domain shift occurs only once and does not reappear. The sequence progresses through distinct target domains in a fixed order $D_{1}\to D_{2}\to\dots\to D_{N}$, with samples from each domain appearing contiguously. The class labels within each domain are independently and randomly selected. The sequence is structured as:
[…]
@@ VIII-A2 Continual Domain Adaptation
VIII-B Further Remarks on Related Works
VIII-B1 Relation to Frequency Domain Learning
Recent advances highlight frequency-based techniques as powerful tools for domain transfer. In domain generalization, frequency analysis has revealed critical insights into model robustness and learning dynamics [50, 51, 52, 53, 54, 55, 56, 57]. For domain adaptation, interpolating image amplitude spectra across styles has proven effective in reducing domain gaps and preventing overfitting to low-level statistics [58, 59, 60, 61]. Motivated by these advancements, we discover that frequency information inherently captures domain characteristics, making it a valuable medium for decoupling mixed target domains – an aspect largely unexplored in prior work. On this basis, we propose FreDA, addressing TTA under mixed domain shifts in Fourier space, with a decentralized adaptation and perturbation mechanism.
VIII-B2 Relation to Multi-Target Domain Adaptation
TTA under mixed distribution shifts shares similarities with multi-target unsupervised domain adaptation (MT-UDA) [62, 11, 12, 13, 63, 64], yet diverges critically in complexity: while MT-UDA assumes static, predefined target domains and leverages labeled source data for explicit domain alignment, TTA operates with no access to source data and must adapt to dynamic, unpredictable target streams in an online manner. This eliminates direct source-target discrepancy computation and demand robust incremental adaptation rather than offline multi-domain optimization. These differences highlight the need for novel methodologies tailored specifically to TTA, going beyond the solutions developed for MT-UDA.
Figure 8: Examples from ImageNet-C under image corruptions. The images showcase a range of corruption types (e.g., noise, blur, and weather distortions) at varying severity levels.
@@ VIII-B2 Relation to Multi-Target Domain Adaptation
VIII-B3 Relation to Decentralized, Federated, Distributed Learning
This work also intersects with decentralized, federated, and distributed learning due to splitting data batches into disjoint subsets and applying decentralized model adaptation. First, while decentralized learning focuses on non-i.i.d. data that is naturally distributed across multiple nodes [65], our approach starts with a centralized batch of target samples. By proactively splitting this data into disjoint subsets, we expose latent non-i.i.d. characteristics, enabling the effective use of decentralized learning techniques. Second, federated learning considers data privacy and model collaborations within decentralized learning [66]. In our case, as target samples are mixed in a batch, data privacy is not a concern. However, similar to federated learning, our approach also involves weight aggregation over subnetworks to enhance their base models. Third, distributed learning aims to improve training efficiency on large-scale datasets by partitioning data for synchronized training [67]. In contrast, our method operates in a real-time fine-tuning context with limited data at one time, hence scalability is less of a concern.
VIII-C Dataset Visualization
@@
IX Conclusion
This paper advances Test-Time Adaptation by tackling the practical challenges posed by heterogeneous data streams. We introduce FreDA, a frequency-based decentralized adaptation framework that partitions incoming samples and performs localized adaptation, while employing distribution-aware augmentation to effectively handle mixed distribution shifts. Empirical results across multiple benchmarks demonstrate the effectiveness of our approach, highlighting its potential to drive future research on adapting to dynamic and diverse domain shifts.
Future Work: FreDA provides a principled framework for adapting to mixed distribution shifts, and we believe several promising directions could further enhance its capabilities. First, beyond the current proactive clustering, future work could investigate adaptive, learnable partitioning strategies that dynamically refine subset boundaries and better capture subtle relationships among target subdomains. Second, our frequency-based augmentation is currently applied at the sample level, which is well-suited for classification. Extending it to patch-level operations in spatial–frequency space may enrich the model’s granularity and open new possibilities for fine-grained tasks such as segmentation. Together, these directions offer natural extensions to broaden FreDA’s applicability in increasingly realistic and dynamic environments.